# Advanced safety filter based on SOS Control Barrier and Lyapunov Functions

### v2

@@ III-B CBF and CLF conditions formalizing the specifications
The next lemma establishes the finite-time convergence property associated with the CLF $V(x)$. Its proof can be found in the appendix.
@@
Given a Lipschitz-continuous controller $u_{s}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ that satisfies condition 17 with $u:=u_{s}(x)$ and renders the safe set $\mathcal{X}_{s}$ forward invariant, then for all trajectories of system 3 starting in the safe set $\mathcal{X}_{s}$ there exists some finite $T\geq 0$ such that $x(t)\in\mathcal{X}_{n}$ for all $t\geq T$.
@@
$\nabla V(x)^{T}\left(f(x)+G(x)u_{n}(x)\right)+d(x)\leq 0.$ (19)
Notice that forward invariance also follows with $d(x)=0$. By adopting a slightly stricter forward invariance condition in 19 involving the strictly positive dissipation rate $d(x)$, we ensure the existence of the safety filter implementation as specified in Section V.
@@
As the volume of the transitional region $\mathcal{X}_{t}$ approaches zero, the proposed safety filter transitions into a safety filter based on a single CBF $V(x)=B_{1}(x)=\ldots=B_{t}(x)$, where the CBF and CLF condition 15 and 17 are automatically satisfied by the forward invariance condition 19 of the legacy controller. Therefore, feasible CBF and CLF candidates can be found (in the limit) if a forward invariant set for the system, governed by the legacy controller, exists and is fully contained within the allowable set.
@@ III-C Problem statement
The challenge of deriving the advanced safety filter can be divided into two subproblems. The first subproblem resolves around finding $B(x)$ and $V(x)$ that fulfill the CBF and CLF condition outlined in the preceding section. This involves constructing a rational controller $u_{\text{SOS}}(x)$ to prove the compatibility of the CLF and CBF conditions in 15 and 17. Furthermore, we necessitate the assurance of the forward invariance condition of the legacy controller as defined in 19.
@@
  • (set containment) the containment conditions in 8 and 10 are satisfied,
  • (finite-time convergence) the CLF condition in 17 is satisfied,
  • (forward-invariance of $\mathcal{X}_{s}$) CBF and CLF conditions in 15 and 17 are satisfied on $\partial\mathcal{X}_{s}$ for the same $u$,
  • (forward-invariance of $\mathcal{X}_{n}$) legacy control condition in 19 is satisfied,
@@
For the second subproblem, a safety filter $u_{s}(x)$ is constructed from $B(x)$ and $V(x)$, which is guaranteed to be feasible and Lipschitz-continuous, entailing a smooth transitioning from the safety filter operation to the legacy control operation.
@@
Given a dynamical system 3, a Lipschitz-continuous legacy controller $u_{n}(x)$, a CBF $B(x)$ and CLF $V(x)$, as specified in Problem 1, find state-feedback control law $u_{s}(x)$ that is feasible for all $x\in\mathbb{R}^{n}$, satisfies the CBF and CLF conditions 15 and 17 with $u:=u_{s}(x)$, coincides with $u_{n}(x)$ on $\mathcal{X}_{n}$ (cf. 9), and is Lipschitz-continuous.
@@
Figure 4: Problem 1 involves searching for compatible CBF $B(x)$ and CLF $V(x)$ using SOS tools. Based on $B(x)$ and $V(x)$, the linear constraints of the QCQP given by $C(x)$, $b(x)$ and $r(x)$ are determined.
@@ IV-A SOS formulation of CBF and CLF conditions
The objective of this subsection is to derive sufficient SOS conditions that encode the CBF condition on $\mathcal{X}_{s}$ and $\mathcal{X}_{n}$, as specified in 15 and 19, and the CLF condition, as specified in 17. We further demonstrate that these SOS conditions are not overly conservative and pose no practical concerns.
The CBF condition in 15 involves the set of active constraints $\text{Act}(x)$, a construction that cannot be directly expressed as an SOS constraint. To address this issue, we first consider the polynomials $B_{i}(x)$, $i\in\mathcal{I}_{t}$, defining the CBF $B(x)$ in 13, as independent CBFs in their own right. By ensuring compatibility among these CBFs $B_{i}(x)$ as proposed in [8], we demonstrate sufficiency of the resulting SOS constraints with respect to CBF condition in 15 (involving the desired CBF $B(x)$). The CBF condition for each $B_{i}(x)$, $i\in\mathcal{I}_{t}$, is expressed as $\forall x\in\{x\in\mathcal{X}_{s}\mid B_{i}(x)=0\},\exists u\in\mathbb{R}^{m}$:
$\nabla B_{i}(x)^{T}\left(f(x)+G(x)u\right)\leq 0.$ (20)
Another challenge in converting conditions 20 and 17 to SOS constraints is the presence of the existential quantifier $\exists u\in\mathbb{R}^{m}$. However, following the approach in [11], we can reformulate these conditions without the need for an existential quantifier. The CBF condition 20 can be expressed as $\forall x\in\{x\in\mathcal{X}_{s}\mid B_{i}(x)=0,\nabla B_{i}(x)^{T}G(x)=0\}$:
$\nabla B_{i}(x)^{T}f(x)\leq 0.$ (21)
Similarly, the CLF condition in 17 can be formulated as $\forall x\in\{x\in\mathcal{X}_{t}\mid\nabla V(x)^{T}G(x)=0\}$:
$\nabla V(x)^{T}f(x)+d(x)\leq 0.$ (22)
In order to render the problem tractable, we restrict our attention to polynomial candidates $B_{i}\in R[x]$, $i\in\mathcal{I}_{t}$, and polynomial candidate $V\in R[x]$. We then apply Putinar’s P-satz, as stated in Theorem 1, on a slightly more stringent conditions compared to 19, 21 and 22, where the inequality is replaced by a strict inequality. These more stringent conditions are summarized as the following empty-set conditions, $i\in\mathcal{I}_{t}$:
$\begin{split}\{x\mid&-\nabla B_{i}(x)^{T}f(x)\leq 0,\nabla B_{i}(x)^{T}G(x)=0,\\ &B_{i}(x)=0,-B_{\mathcal{I}_{t}}(x)\geq 0\}=\emptyset\end{split}$ (23a)
$\begin{split}\{x\mid&-\nabla V(x)^{T}f(x)-d(x)\leq 0,\nabla V(x)^{T}G(x)=0,\\ &V(x)\geq 0,-B_{\mathcal{I}_{t}}(x)\geq 0\}=\emptyset\end{split}$ (23b)
$\begin{split}\{x\mid&\nabla V(x)^{T}\left(f(x)+G(x)u_{n}(x)\right)+d(x)\leq 0,\\ &V(x)=0\}=\emptyset,\end{split}$ (23c)
where $B_{\mathcal{I}_{t}}(x)=\begin{bmatrix}B_{1}(x)&\ldots&B_{t}(x)\end{bmatrix}^{T}$ is a vector of CBFs. The resulting SOS constraints are summarized below. A detailed derivation of the SOS constraint 24 from the empty-set condition 23 is given in the supplementary material in [19].
@@
for some $p_{i}\in(R[x])^{m}$ and $s_{i}\in\Sigma[x]$, $i\in\{0\}\cup\mathcal{I}_{t}$. For clarity, we omit the brackets and variables in the function notation. To achieve the simultaneous satisfaction of CBF and CLF conditions 15 and 17 with a single input $u:=p(x)/s(x)$ on $\partial\mathcal{X}_{s}$, as specified in Problem 1, we equate the polynomials $p_{i}(x)$ and $s_{i}(x)$, $i\in\{0\}\cup\mathcal{I}_{t}$. The resulting SOS constraints are given as:
@@
The polynomials $p(x)$ and $s(x)$ induce (after dividing 25a and 25b by $s(x)$) a rational controller $u_{\text{SOS}}(x):=p(x)/s(x)$, employed to prove compatibility of the CBFs $B_{i}(x)$ and CLF $V(x)$.
The proof of Lemma 2 can be found in the appendix.
@@
If $s(x)>0$ and Assumption 2 is satisfied, then the SOS constraints in 25 are sufficient conditions for the CBF and CLF property in 15, 17 and 19 to hold simultaneously with $u:=p(x)/s(x)$.
The SOS constraints in 25 imply the inequalities since $p\in\Sigma[x]$ implies $p(x)\geq 0$:
$\begin{split}\nabla B_{i}^{T}\left(sf+Gp\right)+\gamma_{0,i}B_{i}-\gamma_{1,i}^{T}B_{\mathcal{I}_{t}}&\leq 0\end{split}$ (26a)
$\begin{split}\nabla V^{T}\left(sf+Gp\right)+sd+\gamma_{0,0}V-\gamma_{1,0}^{T}B_{\mathcal{I}_{t}}&\leq 0\end{split}$ (26b)
$\begin{split}\nabla V^{T}\left(f+Gu_{n}\right)+d+\gamma_{n}V&\leq 0,\end{split}$ (26c)
for some $\gamma_{0,i},\gamma_{n}\in R[x]$, $p\in(R[x])^{m}$, $s,\gamma_{0,0}\in\Sigma[x]$, and $\gamma_{1,i},\gamma_{1,0}\in(\Sigma[x])^{t}$, $i\in\mathcal{I}_{t}$.
Lemma 2 only provides sufficient conditions. However, adding an arbitrary small perturbation $\epsilon>0$ to the SOS constraints in 25 results in necessary conditions when restricting to rational controllers $u_{\text{SOS}}(x)=p(x)/s(x)$. This is because 15, 17 and 19 with $u:=u_{\text{SOS}}(x)$ imply strict inequalities involving $\epsilon$ as follows: $p(x)\leq 0$ implies $p(x)-\epsilon<0$. These strict inequalities imply the SOS constraints 25 perturbed by $\epsilon$ according to Theorem 1. Hence, for all practical purposes, the solution set of the SOS constraints in 25 encompasses the desired set of CBFs and CLFs.
@@ IV-B SOS formulation of input constraints
where $a_{i}$ is the $i$th row of $A$, $i\in\mathcal{I}_{l}$, and $p(x)$ and $s(x)$ are as in 25. Next, we consider the quadratic input constraint
$\mathcal{U}=\{u\mid(u-u_{0})^{T}Q_{u}(u-u_{0})\leq u_{\text{max}}^{2}\},$ (29)
for some $Q_{u}\succeq 0$ and $u_{0}\in\mathbb{R}^{m}$. The quadratic input constraint 29 can be encoded using the SOS constraint involving variables $x\in\mathbb{R}^{n}$ and $v\in\mathbb{R}^{m}$:
$\begin{split}&-v^{T}Q_{u}\left(p(x)-s(x)u_{0}\right)+s(x)u_{\text{max}}^{2}\in\textbf{M}(-f_{u},-B_{\mathcal{I}_{t}}),\end{split}$ (30)
where $f_{u}(x,v):=v^{T}Q_{u}v-u_{\text{max}}^{2}\in R[x,v]$ encodes the input constraint set in 29. The following lemma establishes the implication of the input constraints based on the SOS constraints. The proof is reported in the appendix.
@@
If $s(x)>0$, a solution $p(x)$ and $s(x)$ to the SOS constraint 28, resp. 30, also satisfies the input constraints 27, resp. 29, with $p(x)/s(x)\in\mathcal{U}$ for all $x\in\mathcal{X}_{s}$.
@@ IV-C Safe set and nominal region requirement
In this section, we express the containment conditions of the allowable set $\mathcal{X}_{a}$, the safe set $\mathcal{X}_{s}$ and the nominal region $\mathcal{X}_{n}$, as defined in 14 and 10 (cf. Figure 1), in terms of SOS constraints. A slightly more stringent condition compared to 14, can be expressed as the empty-set condition for all $i\in\mathcal{I}_{t}$:
@@
Using Putinar’s P-satz in Theorem 1, we translate 31 into the SOS constraints for all $i\in\mathcal{I}_{t}$:
@@
Similarly, the containment condition of $\mathcal{X}_{n}$ inside $\mathcal{X}_{s}$ in 10, can be expressed as the empty-set condition for all $i\in\mathcal{I}_{t}$:
@@
Using Putinar’s P-satz, we translate 33 into the SOS constraints for all $i\in\mathcal{I}_{t}$:
@@ IV-D Resulting SOS problem
The solution to Problem 1, i.e., finding suitable functions $B_{i}(x)$, $i\in\mathcal{I}_{t}$, and $V(x)$ can be obtained by solving the following SOS optimization problem:
@@
where 25 encode the CBF and CLF condition, 28 and 30 encode linear and quadratic input constraints, 32 and 34 encode the set containment properties, and a constant $\epsilon>0$ assures that $s(x)>0$.

### v3

@@ III-B CBF and CLF conditions formalizing the specifications
The following lemma establishes the finite-time convergence property induced by the CLF $V(x)$. A detailed proof is provided in Section VIII-A .
@@
Given a Lipschitz-continuous controller $u_{s}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ that satisfies condition 18 with $u:=u_{s}(x)$ and renders the safe set $\mathcal{X}_{s}$ forward invariant, then for all trajectories of system $\dot{x}=f(x)+G(x)u_{s}(x)$ starting in the safe set $\mathcal{X}_{s}$ there exists some finite $T\geq 0$ such that $x(t)\in\mathcal{X}_{n}$ for all $t\geq T$.
@@
$\nabla V(x)^{\top}\left(f(x)+G(x)u_{n}(x)\right)+d(x)\leq 0.$ (20)
Notice that forward invariance also follows with $d(x)=0$. By adopting a slightly stricter forward invariance condition in 20 involving the strictly positive dissipation rate $d(x)$, we are able to proof the existence of a safety filter implementation in Section V.
@@
Compared to approaches that simultaneously establish safety and stability [17], our proposed specifications are less restrictive. In fact, feasible CBF and CLF candidates that satisfy these specifications can always be found, provided that a forward invariant set under the legacy controller exists – including a dissipation term as in 20 – that is fully contained within the allowable set. To such cases, a trivial but valid construction would be to set all barrier and Lyapunov functions equal, such that CBF and CLF condition 16 and 18 are automatically satisfied by the forward invariance condition of the legacy controller.
@@ III-C Problem statement
The advanced safety filter can be divided into two subproblems. The first resolves around finding $B(x)$ and $V(x)$ that fulfill the CBF and CLF condition outlined in the preceding section.
@@
  • (set containment) the containment conditions in 9 and 11 are satisfied,
  • (finite-time convergence) the CLF condition in 18 is satisfied,
  • (forward-invariance of $\mathcal{X}_{s}$) CBF and CLF conditions in 16 and 18 are satisfied on $\partial\mathcal{X}_{s}$ for the same $u$,
  • (forward-invariance of $\mathcal{X}_{n}$) legacy control condition in 20 is satisfied,
@@
For the second subproblem, a safety filter $u_{s}(x)$ is constructed based on the CBF and CLF.
@@
Given a dynamical system 3, a Lipschitz-continuous legacy controller $u_{n}(x)$, a CBF $B(x)$ and CLF $V(x)$, as specified in Problem 1, find state-feedback control law $u_{s}(x)$ that:
  • is feasible for all $x\in\mathbb{R}^{n}$,
  • satisfies the CBF and CLF conditions 16 and 18 with $u:=u_{s}(x)$,
  • coincides with $u_{n}(x)$ on $\mathcal{X}_{n}$ (cf. 10),
  • is Lipschitz-continuous.
@@
Figure 4: Problem 1 involves searching for compatible CBF $B(x)$ and CLF $V(x)$ using SOS tools. Based on $B(x)$ and $V(x)$, the linear constraints of the QP given by $C(x)$, $b(x)$ and $r(x)$ are determined.
@@ IV-A SOS formulation of CBF and CLF conditions
The objective of this subsection is to derive sufficient SOS conditions that encode the CBF condition on $\mathcal{X}_{s}$ and $\mathcal{X}_{n}$, as specified in 16 and 20, and the CLF condition, as specified in 18. We further demonstrate that these SOS conditions are not overly conservative and pose no practical concerns.
The CBF condition in 16 involves the set of active constraints $\text{Act}(x)$, a construction that cannot be directly expressed as an SOS constraint. To address this issue, we first consider the polynomials $B_{i}(x)$, $i\in\mathcal{I}_{t}$, defining the CBF $B(x)$ in 14, as independent CBFs in their own right. By ensuring compatibility among these CBFs $B_{i}(x)$ as proposed in [18], we demonstrate sufficiency of the resulting SOS constraints with respect to CBF condition in 16 (involving the desired CBF $B(x)$). The CBF condition for each $B_{i}(x)$, $i\in\mathcal{I}_{t}$, is expressed as $\forall x\in\{x\in\mathcal{X}_{s}\mid B_{i}(x)=0\},\exists u\in\mathbb{R}^{m}$:
$\nabla B_{i}(x)^{\top}\left(f(x)+G(x)u\right)\leq 0.$ (21)
Another challenge in converting conditions 21 and 18 to SOS constraints is the presence of the existential quantifier $\exists u\in\mathbb{R}^{m}$. However, following the approach in [19], we can reformulate these conditions without the need for an existential quantifier. The CBF condition 21 can be expressed as $\forall x\in\{x\in\mathcal{X}_{s}\mid B_{i}(x)=0,\nabla B_{i}(x)^{\top}G(x)=0\}$:
$\nabla B_{i}(x)^{\top}f(x)\leq 0.$ (22)
Similarly, the CLF condition in 18 can be formulated as $\forall x\in\{x\in\mathcal{X}_{t}\mid\nabla V(x)^{\top}G(x)=0\}$:
$\nabla V(x)^{\top}f(x)+d(x)\leq 0.$ (23)
In order to render the problem tractable, we restrict our attention to polynomial candidates $B_{i}\in R[x]$, $i\in\mathcal{I}_{t}$, and polynomial candidate $V\in R[x]$. We then apply Putinar’s P-satz, as stated in Theorem 1, on a slightly more stringent conditions compared to 20, 22 and 23, where the inequality is replaced by a strict inequality. These more stringent conditions are summarized as the following empty-set conditions, $i\in\mathcal{I}_{t}$:
$\begin{split}\{x\mid&-\nabla B_{i}(x)^{\top}f(x)\leq 0,\nabla B_{i}(x)^{\top}G(x)=0,\\ &B_{i}(x)=0,-B_{\mathcal{I}_{t}}(x)\geq 0\}=\emptyset\end{split}$ (24a)
$\begin{split}\{x\mid&-\nabla V(x)^{\top}f(x)-d(x)\leq 0,\nabla V(x)^{\top}G(x)=0,\\ &V(x)\geq 0,-B_{\mathcal{I}_{t}}(x)\geq 0\}=\emptyset\end{split}$ (24b)
$\begin{split}\{x\mid&\nabla V(x)^{\top}\left(f(x)+G(x)u_{n}(x)\right)+d(x)\leq 0,\\ &V(x)=0\}=\emptyset,\end{split}$ (24c)
where $B_{\mathcal{I}_{t}}(x)=\begin{bmatrix}B_{1}(x)&\ldots&B_{t}(x)\end{bmatrix}^{\top}$ is a vector of CBFs. The resulting SOS constraints are summarized below. A detailed derivation of the SOS constraint 25 from the empty-set condition 24 is given in Section VIII-B.
@@
for some $p_{i}\in(R[x])^{m}$ and $s_{i}\in\Sigma[x]$, $i\in\{0\}\cup\mathcal{I}_{t}$. For clarity, we omit the brackets and variables in the function notation. To achieve the simultaneous satisfaction of CBF and CLF conditions 16 and 18 with a single input $u:=p(x)/s(x)$ on $\partial\mathcal{X}_{s}$, as specified in Problem 1, we equate the polynomials $p_{i}(x)$ and $s_{i}(x)$, $i\in\{0\}\cup\mathcal{I}_{t}$. The resulting SOS constraints are given as:
@@
The polynomials $p(x)$ and $s(x)$ induce (after dividing 26a and 26b by $s(x)$) a rational controller $u_{\text{SOS}}(x):=p(x)/s(x)$, employed to prove compatibility of the CBFs $B_{i}(x)$ and CLF $V(x)$.
The proof of Lemma 2 can be found in Section VIII-C.
@@
If $s(x)>0$ and Assumption 2 is satisfied, then the SOS constraints in 26 are sufficient conditions for the CBF and CLF property in 16, 18 and 20 to hold simultaneously with $u:=p(x)/s(x)$.
The SOS constraints in 26 imply the inequalities since $p\in\Sigma[x]$ implies $p(x)\geq 0$:
$\begin{split}\nabla B_{i}^{\top}\left(sf+Gp\right)+\gamma_{0,i}B_{i}-\gamma_{1,i}^{\top}B_{\mathcal{I}_{t}}&\leq 0\end{split}$ (27a)
$\begin{split}\nabla V^{\top}\left(sf+Gp\right)+sd+\gamma_{0,0}V-\gamma_{1,0}^{\top}B_{\mathcal{I}_{t}}&\leq 0\end{split}$ (27b)
$\begin{split}\nabla V^{\top}\left(f+Gu_{n}\right)+d+\gamma_{n}V&\leq 0,\end{split}$ (27c)
for some $\gamma_{0,i},\gamma_{n}\in R[x]$, $p\in(R[x])^{m}$, $s,\gamma_{0,0}\in\Sigma[x]$, and $\gamma_{1,i},\gamma_{1,0}\in(\Sigma[x])^{\top}$, $i\in\mathcal{I}_{t}$.
Lemma 2 only provides sufficient conditions. However, adding an arbitrary small perturbation $\epsilon>0$ to the SOS constraints in 26 results in necessary conditions when restricting to rational controllers $u_{\text{SOS}}(x)=p(x)/s(x)$. This is because 16, 18 and 20 with $u:=u_{\text{SOS}}(x)$ imply strict inequalities involving $\epsilon$ as follows: $p(x)\leq 0$ implies $p(x)-\epsilon<0$. These strict inequalities imply the SOS constraints 26 perturbed by $\epsilon$ according to Theorem 1. Hence, for all practical purposes, the solution set of the SOS constraints in 26 encompasses the desired set of CBFs and CLFs.
@@ IV-B SOS formulation of input constraints
where $a_{i}$ is the $i$th row of $A$, $i\in\mathcal{I}_{l}$, and $p(x)$ and $s(x)$ are as in 26. Next, we consider the quadratic input constraint
$\mathcal{U}=\{u\mid(u-u_{0})^{\top}Q_{u}(u-u_{0})\leq u_{\text{max}}^{2}\},$ (30)
for some $Q_{u}\succeq 0$ and $u_{0}\in\mathbb{R}^{m}$. The quadratic input constraint 30 can be encoded using the SOS constraint involving variables $x\in\mathbb{R}^{n}$ and $v\in\mathbb{R}^{m}$:
$\begin{split}&-v^{\top}Q_{u}\left(p(x)-s(x)u_{0}\right)+s(x)u_{\text{max}}^{2}\in\textbf{M}(-f_{u},-B_{\mathcal{I}_{t}}),\end{split}$ (31)
where $f_{u}(x,v):=v^{\top}Q_{u}v-u_{\text{max}}^{2}\in R[x,v]$ encodes the input constraint set in 30. The following lemma establishes the implication of the input constraints based on the SOS constraints. The proof is given in Section VIII-D.
@@
If $s(x)>0$, a solution $p(x)$ and $s(x)$ to the SOS constraint 29, resp. 31, also satisfies the input constraints 28, resp. 30, with $p(x)/s(x)\in\mathcal{U}$ for all $x\in\mathcal{X}_{s}$.
@@ IV-C Safe set and nominal region requirement
In this section, we express the containment conditions of the allowable set $\mathcal{X}_{a}$, the safe set $\mathcal{X}_{s}$ and the nominal region $\mathcal{X}_{n}$, as defined in 15 and 11 (cf. Figure 1), in terms of SOS constraints. A slightly more stringent condition compared to 15, can be expressed as the empty-set condition for all $i\in\mathcal{I}_{t}$:
@@
Using Putinar’s P-satz in Theorem 1, we translate 32 into the SOS constraints for all $i\in\mathcal{I}_{t}$:
@@
Similarly, the containment condition of $\mathcal{X}_{n}$ inside $\mathcal{X}_{s}$ in 11, can be expressed as the empty-set condition for all $i\in\mathcal{I}_{t}$:
@@
Using Putinar’s P-satz, we translate 34 into the SOS constraints for all $i\in\mathcal{I}_{t}$:
@@ IV-D Resulting SOS problem
The solution to Problem 1 – that is the synthesis of suitable functions $B_{i}(x)$, $i\in\mathcal{I}_{n_{\!B}}$, and $V(x)$ – can be obtained by solving the following SOS optimization problem:
@@
Here, the constraints encode set the CBF and CLF conditions, the containment properties 9 and 11, and a constant $\epsilon>0$ assures that $s(x)>0$. Additional SOS constraints can be incorporated to capture linear and quadratic input constraint 29 and 31.